OpenStudy (anonymous):

graph each function; identify the domain and range; and compare the graph with the graph of y=1/x. y=1/x+3 +3 domain: Range: Compare:

OpenStudy (anonymous):

i am going to guess that this is \[y=\frac{1}{x+3}+3\]

OpenStudy (anonymous):

compared to the graph of \(y=\frac{1}{x}\) the graph of \(y=\frac{1}{x+3}\) is moved to the LEFT 3 units

OpenStudy (anonymous):

then when you add 3 out at the end to get \[y=\frac{1}{x+3}+3\] it is now UP 3 units net effect is "left 3, up 3"

OpenStudy (anonymous):

domain: makes sure the denominator is not zero, i.e. solve \(x+3=0\) and get \(x=-3\) in one step, so domain is all numbers except \(- 3\)

OpenStudy (anonymous):

as for the range, a fraction is only zero if the numerator is zero so \(\frac{1}{x+3}\) can never be zero since the numerator is \(1\) and therefore \(\frac{1}{x+3}+3\) can never be \(3\) no matter what \(x\) is range is all numbers except \(y=3\)

OpenStudy (anonymous):

how would i graph this